Wolf Creek Pass Tanker

All POL/SITREP's for this site Wolf Creek Pass Tanker
South Fork, CO - EPA Region VIII
Printer Friendly  |   PDF
Wolf Creek Pass Tanker - Removal Polrep

EPA Emergency Response

Region VIII

Subject: POLREP #15
Wolf Creek Pass Tanker
South Fork, CO
Latitude: 37.4748581 Longitude: -106.7816162

From: Pete Stevenson, OSC
Date: 6/3/2015
Reporting Period:

1. Introduction
  1.1 Background
Site Number: Z8D7    Contract Number:  
D.O. Number:      Action Memo Date:  
Response Authority: OPA    Response Type: PRP Oversight
Response Lead: PRP    Incident Category: Removal Action
NPL Status: Non NPL    Operable Unit:
Mobilization Date:      Start Date: 12/4/2009
Demob Date:      Completion Date:  
ERNS No.:    State Notification:
FPN#: E10802    Reimbursable Account #:

1.1.1 Incident Category

1.1.2 Site Description

The tanker accident occurred 11/14/09 at milepost 179, US Hwy 160, Rio Grande County, (approximately 4 ½ miles south of South Fork, Colorado.  An estimated 3,800 gallons of gasoline drained from the overturned tanker across and under the highway into wetlands between the highway and the South Fork Rio Grande River. Location Description of Threat

Groundwater contamination and, at a minimum, dissolved BTEX are impacting the river.  The product continues to bleed out of the bedrock below the highway into a wetlands immediately adjacent to the river.  The river is thought to be habitat for the Rio Grande Cutthroat Trout.  There are many water users downstream.  The wetlands is on property administered by the US Forest Service, Rio Grande National Forest.

 1.1.3 Preliminary Removal Assessment/Removal Site Inspection Results

2. Current Activities
  2.1 Operations Section
    2.1.1 Narrative
11/18: PRP-contracted cleanup contractor arrived on Site for initial recon, followed by crew mobilization. 
11/19: Cleanup contractor initiated detailed surface and sub-surface survey(s) in the wetlands beween the highway and the river, and along riverbanks downstream of the accident site. 
12/01: Cleanup contractor initiated installation of monitoring/recovery/extraction wells and trenches.
12/03: Cleanup contractor initiated recovery of free-product from various wells and trenches via bailing, pumping and sorbents. 
12/21: Contractors temporarily ceased on-site activities due to inclement weather and holiday road restrictions. 
Cleanup contractor notified down-stream water users of activities via posted flyers. 
01/08/10: Interceptor trench completed, including liner installation and insulating blanket coverage. 
01/11: START collected co-located water samples with PRP cleanup contractor. Analytical results are reasonably consistent with PRP sample data.
01/14:  Additional monitoring trenches installed.  One additional recovery hole dug.
02/01:  3-4' of additional snow on-site; many monitoring wells iced up.
02/02:   PRP contractor recovered approximately 1,500 gals of product/water mixture into a vac truck.  Another recovery hole was excavated, which by mid-day was producing gas/water with a very strong gas odor.  Downstream boom frozen in.

March 2010

PRP cleanup contractor(s) continue to recover free product from the various wells and trenches, constructed in the wetlands, between the spill site and the river channel. The cleanup contractor reports measuring up to five inches (5”) of free product in various monitoring wells as well as elevated PID readings at various monitoring stations.  Dissolved phase product continues to be detected in the river.  Due to ice cover, it cannot be determined if a sheen or product is present. 

April 2010

Contaminated water recovery continues.  A meeting was held with interested parties on 4/13/10 to discuss issues related to spring runoff and high water potentially inundating the Site.  The OSC suggested sandbagging the Site to protect it from the river as a first step.  Clean upcontractor collected soil gas data for bioremediation proposal.

April 2011

After another winter, the gasoline continues to bleed out of the bedrock.  The Site appeared abandoned, with trash all around.  The interceptor trench needs to be maintained, the sand bag berm needs to be rebuilt, and contaminated water needs to be pumped out of the recovery holes weekly.  Sampling has to be completed bi-weekly, and the PRP contractor needs to provide weekly reports - the last one received was dated 3/30/11.  Photographs of Site conditions were taken.  USFS personnel are considering revoking the special use permit.  EPA is considering issuing a violation notice pursuant to the cleanup order.  EPA is also considering federalization of the Site.  At a minimum, START will put together a design for an automated pump and treat system to be installed this spring or summer.  PRP contractors are scrambling to comply after receipt of an email from OSC.

December 2011

PRP consultant and cleanup contractor conducted pumping of contaminated water through the summer, approximately 1,000 gallons per week.  Bi-weekly water samples were also collected, the results of which indicate that contaminant levels are decreasing.  MW-1 which is on the Hwy 160 shoulder, was found to contain water and is now being sampled monthly during the winter sampling schedule.  Contaminant levels remain relatively high.  On December 16, 2011, one additional monitoring well was installed on the Hwy 160 shoulder.  The other planned well has been installed.

May 2012
Bioremediation plan to include use of oxygen has been submitted by PRP consultant, and reviewed by all parties, with comments sent back to PRP.  A meeting was conducted with PRP consultants, CDPHE, CDOT, START, and the OSC to discuss comments on May 16, 2012. Plan approved

July 2012.  Bioremediation Pilot Project delayed by newly installed wells being slow to develop.

August 2012.  PRP consultant submitted plan for Pilot Project for oxidation of contaminant with hydrogen peroxide.  RRT consulted.  Determined that H2O2 would break down before it reached the South Fork of the Rio Grande.  Pilot Project approval based on modified quantity of oxidant and total metals sampling and analysis.  Project occurred the first week of September.  Slight reduction in contaminant concentrations seen.  Metals being monitored in groundwater to see if any are mobilized by the H2O2.  A second H2O2 injection occurred in November 2012.

The sand bag berm, interceptor trench, and recovery sumps and trenches were all removed, and the surface was graded in April 2013.  The gasoline odors in the recovery sumps and trenches had decreased significantly, and the April sampling results confirm the continued degradation of the product on-site.

The bioremediation pilot test using oxygen has concluded, and a full scale system has been installed and is operational.  As of February 2014, and based on the quarterly sampling results, the system, and natural attenuation, has decreased the contaminant plume to a much smaller area in the wetlands.  The contamination levels under Hwy. 160 also continue to decrease.

The system was shut down in October of 2014 and restarted in May of 2015.  It is expected to run through the summer and fall when it will be shut down again.  The RP will conduct quarterly sampling with OSC oversight.

2.1.2 Response Actions to Date

See above

2.1.3 Enforcement Activities, Identity of Potentially Responsible Parties (PRPs)

Administrative Order pursuant to Section 311(c) of the Clean Water Act issued to J C Hunt Trucking on April 19, 2010

2.1.4 Progress Metrics

Waste Stream Medium Quantity Manifest # Treatment Disposal
 gas/water       500 gal  001334768  Arvada TreatmentCtr LLC  
 gas/water    1,500 gal  001334763      "              "  
 gas/water  1,500 gal  001334765      "              "
 gas/water  many frac   tanks  NA Air stripping Alamosa WWTP

  2.2 Planning Section
    2.2.1 Anticipated Activities
• PRP will continue coordinating plans and activities with representatives of CDOT, CSP, USFS, BLM and EPA/8 OSC. 
• Bioremediation utilizing oxygen is in place and functioning. 
• Cleanup contractor(s) will perform quarterly sampling of groundwater monitoring wells and one stream sample.  The next sampling is scheduled for August 2015, and occasional site visits will be conducted by the OSC. Planned Response Activities Next Steps
EPA/8 OSC will continue oversight of PRP-lead cleanup activities pursuant to the 311(c) Order.  Oversight of the Bioremediation Project will continue.

2.2.2 Issues

  2.3 Logistics Section
    No information available at this time.

  2.4 Finance Section
    No information available at this time.

  2.5 Other Command Staff
    2.5.1 Safety Officer

2.6 Liaison Officer

2.7 Information Officer

2.7.1 Public Information Officer

2.7.2 Community Involvement Coordinator
 A fact sheet was produced and distributed to the local newspaper after a citizen of South Fork inquired as to the status of the cleanup.  Questions were answered from a local reporter regarding the status of the clean up in April 2012

3. Participating Entities
  No information available at this time.

4. Personnel On Site
  START has been demobilized from the site in mid-2013 at the award of the new START contract.  The OSC is conducting oversight of site activities.

5. Definition of Terms
  No information available at this time.

6. Additional sources of information
  No information available at this time.

7. Situational Reference Materials
  No information available at this time.